OpenStudy (anonymous):

Which number equals the square root of 1, 875.? 5the square root of 15. 5the square root of 3. 25the square root of 3. 15the square root of 15.

OpenStudy (anonymous):

help please

OpenStudy (raden):

setting 1,875 = 625 * 3 so, sqrt(1,875) = sqrt(625 * 3) = sqrt(625) * sqrt(3) can u evaluate the sqrt(625) = ... ?

OpenStudy (anonymous):

25

OpenStudy (raden):

yup, so the final answer is ...

OpenStudy (anonymous):

25 because this is the root of 625 and 625 times 3 is 1875 so the anser is c

OpenStudy (raden):

sqrt(625) * sqrt(3) = 25 sqrt(3) you are right

OpenStudy (anonymous):

thanks you rock